OpenStudy (anonymous):

Write an equation of the line that passes through the point (9,3) and has the slope of 4.

OpenStudy (anonymous):

y-3=4(x-9) y-3=4x-36 y=4x-36+3 y=4x-33

OpenStudy (anonymous):

thank you!!